OpenStudy (anonymous):

Write an equation of line passes through (3,-2) and(5,-2) in standard form

OpenStudy (lgbasallote):

y +2 = -5/7(x-3) 7y+14 = -5x + 15 5x + 7y -1 = 0

OpenStudy (anonymous):

why would it not just be y=-2 since it is a straight line on the negative 2 with no slope

OpenStudy (lgbasallote):

y - y1 + (y2-y1/x2-x1)(x-x1)

OpenStudy (anonymous):

what?

OpenStudy (lgbasallote):

it's the two point form...y - y sub 2 = slope (x - x sub 1) slope = y sub 2 - y sub 1 all over x sub 2 - x sub 1 x sub 1 means the x of your first coordinate y sub 1 means the y of your first coordinate x sub 2 means the x of your second coordinate y sub 2 means the y of your second coordinate did you get all that?

OpenStudy (anonymous):

yes but i still get a slope of 0